You are designing some shielding for an x-ray machine. A single 1.0 mm thick sheet of steel blocks 25% of all incident x-rays and allows 75% through.

How thick does shielding made out of this steel need to be to block 90% of all incident x-rays and allow only 10% through? (to 2.s.f and in mm)
1
Expert's answer
2020-05-25T11:16:26-0400

We know that the amount radiation that passes thru decreases inversely with thickness squared:


"I=I_0\frac{k}{t^2}\rightarrow\frac{I}{I_0}=\frac{k}{t^2}."


The proportionality coefficient "k" is the same for all cases, therefore


"k=\frac{I}{I_0}t^2."

The value of the coefficient is


"k=0.75\cdot1^2=0.75."

10% of the radiation passed corresponds, therefore, to a thickness of


"t=\sqrt{k\frac{I_0}{I}}=\sqrt{0.75\frac{1}{0.1}}=2.7\text{ mm}."